OpenStudy (anonymous):

What is the molar mass of 1-butene if 5.38 × 10^16 molecules of 1-butene weigh 5.00 μg?

OpenStudy (aaronq):

Make a ratio: \(\sf \dfrac{mass~of~sample}{\#~of~molecules}=\dfrac{Molar~mass}{Avogadro's~\#}\)

OpenStudy (anonymous):

@aaronq Is it okey if you explain it by number? :)

OpenStudy (aaronq):

explain by number? what do you mean?

OpenStudy (anonymous):

@aaronq Sorry English not my native language I mean like that 5.00/5.38*10^1......

OpenStudy (aaronq):

so 5.00 \(\mu g\)/(5.38 × 10^16 molecules)=molar mass/(6.022*10^23)

OpenStudy (anonymous):

@aaronq Thanks :)

OpenStudy (aaronq):

no problem